# Impact of flavour coupling on $SO(10)$-inspired leptogenesis

### v1

@@ Abstract
[…]
@@ 1 Introduction
We are entering an important stage for the prospect of testing high scale thermal leptogenesis [1] within minimal type I seesaw models [2], the minimal scenario of leptogenesis [3]. Colliders have found no evidence of new physics at the TeV scale and below so far, placing strong constraints (and doubts) on low scale leptogenesis scenarios. At the same time, low energy neutrino experiments continue to progress steadily and in the next years they will be able to provide information on the neutrino mixing unknowns and to test the absolute neutrino mass scale in regions of the parameter space that are quite important for high scale leptogenesis scenarios. Moreover, the discovery of gravitational waves [4] is stimulating many new ideas on how to test high scale leptogenesis [5, 6, 7, 8]. For these reasons, the often proclaimed statement for which high scale leptogenesis scenarios are untestable is today outdated.
@@
An example of well motivated, testable, high scale leptogenesis scenario is SO10INLEP. This relies on quite minimal assumptions on the Dirac neutrino mass matrix, so-called $SO(10)$-inspired conditions [9]. They are typically realised within various grandunified models, in particular within $SO(10)$ models. With these assumptions, and barring the highly fine-tuned compact spectrum case [10, 11, 12], the emerging spectrum of RH neutrinos is strongly hierarchical. In particular the lightest RH neutrino, $N_{1}$, has, typically, a mass $M_{1}\sim(10^{5}$$10^{6})\,{\rm GeV}$, well below the lower bound for successful leptogenesis [13, 14]. On the other hand, the next-to-lightest RH neutrino, $N_{2}$, turns out to have a mass with just the correct order-of-magnitude, $M_{2}\sim 10^{11}\,{\rm GeV}$, to produce an asymmetry able to explain the observed value, realising successful $N_{2}$-leptogenesis [15]. In order for the $C\!P$ asymmetries of $N_{2}$ to be sufficiently large, the existence of a third heavier RH neutrino species, $N_{3}$, is necessary to produce the needed interference in $N_{2}$ decays. This nicely fits within the properties of $SO(10)$-models that indeed, notoriously, predict the existence of three RH neutrino species, one for each of the three family $SO(10)$ fermionic sixteenth-plet representations. When $SO(10)$-inspired conditions are imposed, typically one obtains $M_{3}\sim(10^{14}$$10^{16}$) GeV. However, notice that for successful thermal leptogenesis it is sufficient that the reheat temperature $T_{\rm RH}\gtrsim M_{2}\sim 10^{11}\,{\rm GeV}$, since only the $N_{2}$’s need to thermalise, while $N_{3}$ plays a role at the virtual level only, in the interfering loop diagrams for $N_{2}$-decays. Still this is a crucial role, since this interference is the source of the $N_{2}$-$C\!P$ asymmetries and, ultimately, of the observed baryon asymmetry. For this reason, when $N_{3}$ decouples, in the limit $M_{3}\gg 10^{16}\,{\rm GeV}$, all flavoured $C\!P$ asymmetries of $N_{2}$ tend to vanish. This yields an upper bound on $M_{3}$ and, consequently, from the seesaw formula, a lower bound on the lightest neutrino mass [16]. Although one can have a sizeable asymmetry generated from $N_{2}$-decays at very high temperatures, $T\sim M_{2}$, for long time such a strongly hierarchical RH neutrino spectrum has been regarded as an obstacle to reproduce the correct observed baryon asymmetry within SO10INLEP. The reason is that, within an unflavoured description of SO10INLEP, the washout from $N_{1}$ at $T\sim M_{1}$ is necessarily strong and a successful $N_{2}$-leptogenesis scenario cannot be realised for such a strongly hierarchical spectrum. However, when flavour effects are included [17, 18], the calculation of the $B-L$ asymmetry splits into three different contributions, one for each charged lepton flavour [19]. The strong washout from $N_{1}$ also splits into three weaker components. In this way an allowed region in the space of low energy neutrino parameters opens up and successful SO10INLEP can be attained [16, 20]. The asymmetry is dominantly produced in the tauon flavour. Subdominant muon-dominated flavour solutions also exist, while electron-dominated solutions just fall short of reproducing the observed asymmetry. [^1] They appear marginally in a supersymmetric version [21]. We should also mention that there is a second way to circumvent the $N_{1}$ washout, even when this is strong in all three flavours [21]. If one the ligthest Dirac neutrino mass is at least two orders of magnitude lower than the up quark mass, then $M_{1}$ gets lower than the sphaleron freeze-out temperature $T_{\rm sph}^{\rm off}\simeq 132\,{\rm GeV}$ [22]. In this case the washout acts only on the lepton asymmetry but not on the observed baryon asymmetry.
@@
It is interesting that successful SO10INLEP is realised only within a region of the entire low energy neutrino parameter space. This yields constraints, and predictions, to be confronted with the experimental results [20]. A first important feature of SO10INLEP predictions is that the allowed region for inverted ordered neutrino masses (IO) is marginal [20, 12], since it necessarily requires the atmospheric neutrino mixing angle to be in the second octant and a lower bound on the lightest neutrino mass. For $\theta_{23}\lesssim 50^{\circ}$, as suggested by latest neutrino oscillation data global analyses at $3\sigma$ [23], one has a lower bound on the lightest neutrino mass $m_{1}\gtrsim 20\,{\rm meV}$, corresponding to $\sum_{i}m_{i}\gtrsim 130\,{\rm meV}$. This is in tension with the cosmological upper bound on the sum of neutrino masses obtained combining CMB anisotropy, baryon acoustic oscillation observations and supernovae data [24, 25, 26, 27]
@@
For this reason, at $3\sigma$, SO10INLEP is viable only for normal ordering (NO) and this is in nice agreement with the latest results from atmospheric neutrino oscillation experiments, that tend to favour NO [28]. This is also confirmed by the most recent global analyses disfavouring IO at $\sim 2.5\sigma$ [23]. Therefore, in the following we will focus exclusively on NO and we will always refer to the cosmological upper bound on the lightest neutrino mass
@@
As mentioned, the most interesting constraint from successful SO10INLEP is the existence of the lower bound on the lightest neutrino mass [16, 20]. This depends in a non-trivial way on the $C\!P$-violating Dirac phase $\delta$ and the atmospheric mixing angle $\theta_{23}$ [29]. There is a large region of the plane $\delta$ versus $\theta_{23}$ that is already incompatible with SO10INLEP at 95% C.L. since the cosmological upper bound on neutrino masses Eq. (1) is violated. In the semi-hierarchical regime, for $10\,{\rm meV}\lesssim m_{1}\lesssim 30\,{\rm meV}$, most stringent constraints hold within SO10INLEP. In particular, SO10INLEP is incompatible with the atmospheric mixing angle in the second octant. On the other hand, this is also a very interesting neutrino mass regime, since only for $m_{1}$ in the (10–30) meV range ST-SO10INLEP can be realised [30]. This is a special scenario where, in addition to the condition of successful leptogenesis, one can also wash out a pre-existing asymmetry as large as $10^{-1}$ [31]. For a hierarchical RH neutrino spectrum, without even imposing $SO(10)$-inspired conditions, successful strong thermal leptogenesis can only be realised for quite special conditions: a tauon-dominated $N_{2}$-leptogenesis scenario [31] and for $m_{1}\gtrsim 10\,{\rm meV}$ [32]. It is then quite non trivial that within successful SO10INLEP, one can also satisfy the conditions for strong thermal leptogenesis for a subset of the solutions. In particular, they can only be satisfied for $m_{1}\lesssim 30\,{\rm meV}$, just coinciding with the current cosmological upper bound Eq. (2). This is another important point showing how we are now just entering a crucial stage in testing SO10INLEP and its strong thermal case ST-SO10INLEP.
For this reason, it is of great importance to understand how solid and accurate these theoretical predictions are. The constraints on the low energy neutrino parameters have been derived in various papers and with independent codes. They have also been derived analytically, expressing the RH neutrino mass spectrum and mixing matrix in terms of the low energy neutrino parameters [12, 33]. In this way one obtains an analytical expression of the final baryon asymmetry in terms of the low energy neutrino parameters. However, the calculation of the asymmetry has been done within a set of approximations that neglects different effects. Three of them are at the level of field theory description:
  • Flavour coupling effects from spectator processes [34, 35, 36].
  • Running of parameters from radiative corrections.
  • Partial equilibration of spectator processes [37].
@@
  • Momentum dependence (requiring solution of the full Boltzmann equations) [38].
  • Decoherence (requiring solution of density matrix equation) [34, 17].
[…]
@@
This is the main aim of the paper, that is structured in the following way. In Section 2 we review SO10INLEP, discussing the assumptions and the equations to calculate the asymmetry, describing the results that we obtained ignoring flavour coupling (we have re-derived then once more confirming previous analyses and updating the experimental constraints in the plots) and comparing the predictions to the latest experimental constraints. In Section 3 we discuss flavour coupling and how this modifies the calculation of the final asymmetry. We show the results for some benchmark points and compare them to the case when flavour coupling is ignored. We also point out how the calculation of the asymmetry becomes much more CPU time consuming. For that reason, the derivation of the hyper-surface in the space of parameters, giving the allowed region, becomes much more challenging to derive. This is indeed so convoluted to require at least one million of points for a precise determination. The use of the analytical expressions for the RH neutrino masses and mixing matrix [12, 33] greatly help in this respect. We show the results of the scatter plots giving both the three-dimensional projection of the hypersurface in the space ($\delta,\theta_{23},m_{1}$), as previously done in [29] neglecting flavour coupling, and different two-dimensional projections. We also highlight the different flavour dominance for each point, confirming that, even including flavour coupling effects, the bulk of the solutions, approximately $90\%$, are tauonic solutions, while muonic solutions represent a subdominant $10\%$ contribution. However, when flavour coupling is included, we show that a very small $0.1\%$ contribution of electronic solutions appear and also new muonic solutions. These new muonic solutions fall in a region of the parameter space that would otherwise be excluded. [^3] A special subset of the muonic solutions that we obtain were also found within a specific Pati-Salam grandunified model with discrete flavour symmetries [41]. However, this special subset is now excluded by the latest neutrino oscillation experimental results since they require $\theta_{23}\simeq 54^{\circ}$. In Section 4 we discuss ST-SO10INLEP, showing how the calculation of the evolution of a large pre-existing asymmetry, and its relic value, is modified by the account of flavour coupling. Here, we show how the allowed region not only survives but even, for some values of $\delta$, allows slightly larger values of $\theta_{23}$. Moerover, the allowed range of $m_{1}$ values also slightly englarges. Finally, in Section 5, we draw some final remarks on the importance of (ST-)SO10INLEP and of our results within the current quest for new physics.
@@ 2 $SO(10)$-inspired leptogenesis
In this section we review general features of SO10INLEP without the inclusion of flavour coupling. We briefly discuss the set of assumptions that define SO10INLEP and then we show how the requirement of successful SO10INLEP produces constraints on the low energy neutrino parameters that we compare with the latest neutrino oscillation experimental results. In this way we can discuss the status of SO10INLEP showing that there is currently a non-trivial compatibility between SO10INLEP predictions and low energy neutrino experimental data. We also show how with the new cosmological upper bound on neutrino masses, we are starting to test a very interesting neutrino mass range for SO10INLEP.

### v2

@@ Abstract
We discuss the impact of flavour coupling on the predictions of low energy neutrino parameters from $SO(10)$-inspired leptogenesis (SO10INLEP). The right-handed (RH) neutrino mass spectrum is strongly hierarchical so that successful leptogenesis relies on generating the asymmetry from next-to-lightest RH neutrino decays ($N_{2}$-leptogenesis) and circumventing the lightest RH neutrino washout. These two conditions yield distinctive predictions such as a lower bound on the lightest neutrino mass $m_{1}\gtrsim 1\,{\rm meV}$. We first review the status of SO10INLEP, noticing how cosmological observations are now testing a particular neutrino mass window, $m_{1}\simeq(10$$30)\,{\rm meV}$, where only the first octant is allowed and a large range of values for the Dirac phase is excluded. Including flavour coupling, we find that the lower bound relaxes to $m_{1}\gtrsim 0.65\,{\rm meV}$. Moreover, new muon-dominated solutions appear slightly relaxing the upper bound on the atmospheric mixing angle. We also study the impact on strong thermal SO10INLEP (ST-SO10INLEP) scenario where, in addition to successful leptogenesis, one can washout a large pre-existing asymmetry. Contrarily to naive expectations, for which flavour coupling could jeopardise the scenario allowing a large pre-existing asymmetry to survive unconditionally, we show, and explain analytically, that ST-SO10INLEP is still viable within almost the same allowed region of parameters. There is even a slight relaxation of the $m_{1}$ viable window from (9–30)meV to (4–40)meV for a $10^{-3}$ pre-existing asymmetry. The new results from atmospheric neutrinos, mildly favouring normal ordering and first octant, are now in nice agreement with the predictions of ST-SO10INLEP. Intriguingly, the predicted $0\nu\beta\beta$ signal is starting to be within the reach of KamLAND-Zen.
@@ 1 Introduction
We are entering an important stage for the prospect of testing high scale thermal leptogenesis fy within minimal type I seesaw models seesaw , the minimal scenario of leptogenesis Blanchet:2012bk . Colliders have found no evidence of new physics at the TeV scale and below so far, placing strong constraints (and doubts, certainly from us) on low scale leptogenesis scenarios. At the same time, low energy neutrino experiments continue to progress steadily and in the next years they will be able to provide information on the neutrino mixing unknowns and to test the absolute neutrino mass scale in regions of the parameter space that are quite important for high scale leptogenesis scenarios. Moreover, the discovery of gravitational waves LIGOScientific:2016aoc is stimulating many new ideas on how to test high scale leptogenesis Dror:2019syi , DiBari:2020bvn , DiBari:2021dri , Fu:2022lrn . For these reasons, the often proclaimed statement for which high scale leptogenesis scenarios are untestable is today outdated.
@@
An example of well motivated, testable, high scale leptogenesis scenario is SO10INLEP. This relies on quite minimal assumptions on the Dirac neutrino mass matrix, so-called $SO(10)$-inspired conditions SO10inspired . They are typically realised within various grandunified models, in particular within $SO(10)$ models. With these assumptions, and barring the highly fine-tuned compact spectrum case Akhmedov:2003dg , compact , decrypting , the emerging spectrum of RH neutrinos is strongly hierarchical. In particular the lightest RH neutrino, $N_{1}$, has, typically, a mass $M_{1}\sim(10^{5}$$10^{6})\,{\rm GeV}$, well below the lower bound for successful leptogenesis di , cmb . On the other hand, the next-to-lightest RH neutrino, $N_{2}$, turns out to have a mass with just the correct order-of-magnitude, $M_{2}\sim 10^{11}\,{\rm GeV}$, to produce an asymmetry able to explain the observed value, realising successful $N_{2}$-leptogenesis geometry . In order for the CPC\!P asymmetries of $N_{2}$ to be sufficiently large, the existence of a third heavier RH neutrino species, $N_{3}$, is necessary to produce the needed interference in $N_{2}$ decays. This nicely fits within the properties of $SO(10)$-models that indeed, notoriously, predict the existence of three RH neutrino species, one for each of the three family $SO(10)$ fermionic sixteenth-plet representations. When $SO(10)$-inspired conditions are imposed, typically one obtains $M_{3}\sim(10^{14}$$10^{16}$) GeV. However, for successful thermal leptogenesis it is sufficient that the reheat temperature $T_{\rm RH}\gtrsim M_{2}\sim 10^{11}\,{\rm GeV}$, since only the $N_{2}$’s need to thermalise, while $N_{3}$ plays a role at the virtual level only, in the interfering loop diagrams for $N_{2}$-decays. Still this is a crucial role, considering that this interference is the source of the $N_{2}$-CPC\!P asymmetries and, ultimately, of the observed baryon asymmetry. For this reason, when $N_{3}$ decouples, in the limit $M_{3}\gg 10^{16}\,{\rm GeV}$, all flavoured CPC\!P asymmetries of $N_{2}$ tend to vanish. This yields an upper bound on $M_{3}$ and, consequently, from the seesaw formula, a lower bound on the lightest neutrino mass riotto1 . Although one can have a sizeable asymmetry generated from $N_{2}$-decays at very high temperatures, $T\sim M_{2}$, for long time such a strongly hierarchical RH neutrino spectrum has been regarded as an obstacle to reproduce the correct observed baryon asymmetry within SO10INLEP. The reason is that, within an unflavoured description of SO10INLEP, the washout from $N_{1}$ at $T\sim M_{1}$ is necessarily strong and a successful $N_{2}$-leptogenesis scenario cannot be realised for such a strongly hierarchical spectrum. However, when flavour effects are included flavoureffects , flavourlep , the calculation of the BLB-L asymmetry splits into three different contributions, one for each charged lepton flavour vives . The strong washout from $N_{1}$ also splits into three weaker components. In this way an allowed region in the space of low energy neutrino parameters opens up and successful SO10INLEP can be attained riotto1 , riotto2 . The asymmetry is dominantly produced in the tauon flavour. Subdominant muon-dominated flavour solutions also exist, while electron-dominated solutions just fall short of reproducing the observed asymmetry. [^1] They appear marginally in a supersymmetric version susy . We should also mention that there is a second way to circumvent the $N_{1}$ washout, even when this is strong in all three flavours susy . If one the ligthest Dirac neutrino mass is at least two orders of magnitude lower than the up quark mass, then $M_{1}$ gets lower than the sphaleron freeze-out temperature $T_{\rm sph}^{\rm off}\simeq 132\,{\rm GeV}$ DOnofrio:2014rug . In this case the washout acts only on the lepton asymmetry but not on the observed baryon asymmetry.
@@
It is interesting that successful SO10INLEP is realised only within a region of the entire low energy neutrino parameter space. This yields constraints, and predictions, to be confronted with the experimental results riotto2 . A first important feature of SO10INLEP predictions is that inverted ordered neutrino masses (IO) are very marginally allowed riotto2 , decrypting , since it necessarily requires the atmospheric neutrino mixing angle to be in the second octant and a lower bound on the lightest neutrino mass. For $\theta_{23}\lesssim 50^{\circ}$, as suggested by latest neutrino oscillation data global analyses at $3\sigma$ nufit24 , one has a lower bound on the lightest neutrino mass $m_{1}\gtrsim 20\,{\rm meV}$, corresponding to $\sum_{i}m_{i}\gtrsim 130\,{\rm meV}$. This is in tension with the cosmological upper bound on the sum of neutrino masses obtained combining CMB anisotropy, baryon acoustic oscillation observations and supernovae data Planck:2018vyg , Allali:2024aiv , Naredo-Tuero:2024sgf , DESI:2025ejh
@@
For this reason, at $3\sigma$, SO10INLEP is viable only for normal ordering (NO) and this is in nice agreement with the latest results from atmospheric neutrino oscillation experiments, that tend to favour NO at $92.3\%$ C.L. Super-Kamiokande:2023ahc . This is strengthened by the most recent global analyses disfavouring IO at $\sim 2.5\sigma$ nufit24 . Therefore, in the following, we will focus exclusively on NO and we will always refer to the cosmological upper bound on the lightest neutrino mass
@@
As mentioned, the most interesting constraint from successful SO10INLEP is the existence of the lower bound on the lightest neutrino mass riotto1 , riotto2 . This depends in a non-trivial way on the CPC\!P-violating Dirac phase $\delta$ and the atmospheric mixing angle $\theta_{23}$ DiBari:2020plh . There is a large region of the plane $\delta$ versus $\theta_{23}$ that is already incompatible with SO10INLEP at 95% C.L. since the cosmological upper bound on neutrino masses Eq. (1) is violated. In the semi-hierarchical regime, for $10\,{\rm meV}\lesssim m_{1}\lesssim 30\,{\rm meV}$, most stringent constraints hold within SO10INLEP. In particular, SO10INLEP is incompatible with the atmospheric mixing angle in the second octant. On the other hand, this is also a very interesting neutrino mass regime, since only for $m_{1}$ in the (10–30) meV range ST-SO10INLEP can be realised ST . This is a special scenario where, in addition to the condition of successful leptogenesis, one can also wash out a pre-existing asymmetry as large as $10^{-1}$ problem . For a hierarchical RH neutrino spectrum, without even imposing $SO(10)$-inspired conditions, successful strong thermal leptogenesis can only be realised for quite special conditions: a tauon-dominated $N_{2}$-leptogenesis scenario problem and for $m_{1}\gtrsim 10\,{\rm meV}$ sophie . It is then quite non trivial that within successful SO10INLEP, one can also satisfy the conditions for strong thermal leptogenesis for a subset of the solutions. In particular, they can only be satisfied for $m_{1}\lesssim 30\,{\rm meV}$, just coinciding with the current cosmological upper bound Eq. (2). This is another important point showing how we are now just entering a crucial stage in testing SO10INLEP and its strong thermal case ST-SO10INLEP.
For this reason, it is of great importance to understand how solid and accurate these theoretical predictions are. The constraints on the low energy neutrino parameters have been derived in various papers and with independent codes. They have also been derived analytically, expressing the RH neutrino mass spectrum and mixing matrix in terms of the low energy neutrino parameters decrypting , full . In this way one obtains an analytical expression of the final baryon asymmetry in terms of the low energy neutrino parameters. However, the calculation of the asymmetry has been done within a set of approximations that neglects different effects. Three of them are at the level of field theory description:
@@
In this paper, encouraged by the current agreement of (ST-)SO10INLEP predictions with neutrino oscillation experiment results and by the fact that absolute neutrino mass scale experiments are starting to test a crucial regime, as discussed above, we start implementing these effects, studying the impact of flavour coupling effects on (ST-)SO10INLEP. This is certainly the most urgent extension of the calculation of the asymmetry within (ST-)SO10INLEP to be considered. In particular, in the case of ST-SO10INLEP one can even legitimately suspect that flavour coupling could jeopardise the whole scenario. This is because a contribution from a large pre-existing asymmetry might survive until the present, leaking from one flavour to another, thus escaping both $N_{2}$ and $N_{1}$ washout. As we will discuss in detail, and anticipated in the abstract, this does not happen and, ultimately, even in the ST-SO10INLEP case, the account of flavour coupling yields just some slight corrections that, however, might prove to be important when atmospheric neutrino mixing angle and Dirac phase will be measured with errors $\sim 1^{\circ}$ and $\sim 10^{\circ}$, respectively, by DUNE+T2HK Ballett:2016daj . At the same time we will highlight some new important aspects of (ST-)SO10INLEP predictions. When these are confronted with the latest results from global analyses on neutrino mixing parameters, we show how the current best fit for NO and first octant, nicely agrees with (ST-)SO10INLEP predictions, removing the tension that was existing with previous results chianese favouring second octant for $\theta_{23}$. Moreover, as we mentioned, even without imposing strong thermal leptogenesis, for $m_{1}\sim(10$$30)\,{\rm meV}$ the second octant is not compatible with SO10INLEP constraints. In addition, there are large ranges of $\delta$ that are excluded even for $m_{1}\lesssim 10\,{\rm meV}$. Therefore, during next years, neutrino oscillation experiments will either be able to increase the statistical significance of the agreement or of course, will be able to rule out SO10INLEP. It is then important, not to draw incorrect conclusions, to start a systematic study of theoretical uncertainties to reduce them at the level of the expected experimental errors.
@@
This is the main aim of the paper, that is structured in the following way. In Section 2 we review SO10INLEP, discussing the assumptions and the equations to calculate the asymmetry, describing the results that we obtained ignoring flavour coupling (we have re-derived then once more confirming previous analyses and updating the experimental constraints in the plots) and comparing the predictions to the latest experimental constraints. In Section 3 we discuss flavour coupling and how this modifies the calculation of the final asymmetry. We show the results for some benchmark points and compare them to the case when flavour coupling is ignored. We also point out how the calculation of the asymmetry becomes much more CPU time consuming. For that reason, the derivation of the hyper-surface in the space of parameters, giving the allowed region, becomes much more challenging to derive. This is indeed so convoluted to require at least one million of points for a precise determination. The use of the analytical expressions for the RH neutrino masses and mixing matrix decrypting , full greatly help in this respect. We show the results of the scatter plots giving both the three-dimensional projection of the hypersurface in the space ($\delta,\theta_{23},m_{1}$), as previously done in DiBari:2020plh neglecting flavour coupling, and different two-dimensional projections. We also highlight the different flavour dominance for each point, confirming that, even including flavour coupling effects, the bulk of the solutions, approximately $90\%$, are tauonic solutions, while muonic solutions represent a subdominant $10\%$ contribution. However, when flavour coupling is included, we show that a very small $0.1\%$ contribution of electronic solutions appear and also new muonic solutions. These new muonic solutions fall in a region of the parameter space that would otherwise be excluded. [^3] A special subset of the muonic solutions that we obtain were also found within a specific Pati-Salam grandunified model with discrete flavour symmetries DiBari:2015oca . However, this special subset is now excluded by the latest neutrino oscillation experimental results since they require $\theta_{23}\simeq 54^{\circ}$. In Section 4 we discuss ST-SO10INLEP, showing how the calculation of the evolution of a large pre-existing asymmetry, and its relic value, is modified by the account of flavour coupling. Here, we show how the allowed region not only survives but even, for some values of $\delta$, allows slightly larger values of $\theta_{23}$. Moerover, the allowed range of $m_{1}$ values also slightly englarges. Finally, in Section 5, we draw some final remarks on the importance of (ST-)SO10INLEP and of our results within the current quest for new physics.
@@ 2 $SO(10)$-inspired leptogenesis
In this section we review general features of SO10INLEP, neglecting flavour coupling. We briefly discuss the set of assumptions that define SO10INLEP and then we show how the requirement of successful SO10INLEP produces constraints on the low energy neutrino parameters that we compare with the latest neutrino oscillation experimental results. In this way we can discuss the status of SO10INLEP showing that there is currently a non-trivial compatibility between SO10INLEP predictions and low energy neutrino experimental data. We also show how with the new cosmological upper bound on neutrino masses, we are starting to test a very interesting neutrino mass range for SO10INLEP.